OpenStudy (yourboydusty):

What is the volume of a sphere with radius 5? A. 33.33 PI units3 B. 100 PI units3 C. 126.67 PI units3 D. 166.67 PI units3

OpenStudy (studygurl14):

Vollume of Sphere: \(\huge\frac{4}{3}\Large\pi r^3\)

OpenStudy (yourboydusty):

So the answer is A?

OpenStudy (yourboydusty):

@3mar

OpenStudy (studygurl14):

I am not sure what PI units are...

OpenStudy (yourboydusty):

oh its not PI units it just means you dont use the pi symbol in your equation

OpenStudy (yourboydusty):

@StudyGurl14

OpenStudy (studygurl14):

Oh! then no, not a

OpenStudy (studygurl14):

\(\huge\frac{4}{3}\Large(5)^3\)

OpenStudy (yourboydusty):

how so...when i do the math i get 33.33

OpenStudy (yourboydusty):

5*5=25 4/3*25/3=100/3 100/3=33.33

OpenStudy (yourboydusty):

@StudyGurl14

OpenStudy (studygurl14):

It is 5 times 5 times 5

OpenStudy (yourboydusty):

166.6 then?

OpenStudy (yourboydusty):

so its D?

OpenStudy (yourboydusty):

@StudyGurl14

OpenStudy (studygurl14):

yes

OpenStudy (yourboydusty):

kk thanks

OpenStudy (3mar):

Sorry! I was no here.